An element with mass 820 grams decays by 26.8% per minute. How much of the element is remaining after 18 minutes, to the nearest 10th of a gram?


1
Expert's answer
2022-04-11T13:50:32-0400

Let "m_0 = 820 \space g,"

"\alpha = 26.8\space\% = 0.268,"

"T = 18\space min,"

"m_t" - mass we have after time of decay "t"


We have:

"m_{t + 1} = m_t(1 - \alpha)"

This is a geometric progression with a common ratio equal to "(1 - \alpha)"

Its general term is represented by the formula "m_t = m_0(1 - \alpha)^t"


"m_T = m_0(1 - \alpha)^T =\\
= 820\cdot(1 - 0.268)^{18}\space g = 3.0\space g"